# DP-SPRT: Differentially Private Sequential Probability Ratio Tests

### v1

@@ Abstract
We revisit Wald’s celebrated Sequential Probability Ratio Test for sequential tests of two simple hypotheses, under privacy constraints. We propose DP-SPRT, a wrapper that can be calibrated to achieve desired error probabilities and privacy constraints, addressing a significant gap in previous work. DP-SPRT relies on a private mechanism that processes a sequence of queries and stops after privately determining when the query results fall outside a predefined interval. This OutsideInterval mechanism improves upon naive composition of existing techniques like AboveThreshold, potentially benefiting other sequential algorithms. We prove generic upper bounds on the error and sample complexity of DP-SPRT that can accommodate various noise distributions based on the practitioner’s privacy needs. We exemplify them in two settings: Laplace noise (pure Differential Privacy) and Gaussian noise (Rényi differential privacy). In the former setting, by providing a lower bound on the sample complexity of any $\varepsilon$-DP test with prescribed type I and type II errors, we show that DP-SPRT is near optimal when both errors are small and the two hypotheses are close. Moreover, we conduct an experimental study revealing its good practical performance.
[…]
@@ 1 Introduction
Related Work. Differential privacy (DP) (Dwork et al.,, 2006) has emerged as the standard framework for addressing privacy concerns in statistical analyses. By ensuring that the presence or absence of any single data point cannot significantly affect the outcome, DP offers a principled method to protect individual data while enabling meaningful statistical inference. Researchers have explored DP hypothesis testing– primarily in the fixed-sample setting. For example, Gaboardi et al., (2016) studies private chi-squared testing. Canonne et al., (2019) and Couch et al., (2019) investigate on optimal private tests for simple hypotheses, and nonparametric hypothesis testing with DP, respectively. These approaches, however, are not directly applicable to sequential testing due to the adaptive nature of sample collection.
Wang et al., (2022) explored the privatization of SPRT by adding Laplace noise to thresholds and using the exponential mechanism for the final decision. However, due to their specific use case, their approach provides a weaker notion of privacy that only requires that critical privacy failures do not happen only for a subset of plausible datasets. In contrast, our work maintains DP throughout the testing process while working to minimize stopping time. Zhang et al., (2022) developed PrivSPRT, a private version of the SPRT using Gaussian noise with Renyi DP guarantees (Mironov,, 2017). The authors provide upper bounds on the error probabilities given fixed thresholds and noise level, but do not specify how to calibrate these thresholds to achieve desired Type I and Type II error rates while maintaining a desired privacy guarantee. In practice, this forces to rely on Monte Carlo simulation to determine appropriate thresholds, a process that may itself leak privacy in certain scenarios (Papernot and Steinke,, 2021). In this work, we overcome this limitation and provide an explicit threshold calibration to guarantee specified upper bounds on the errors as well as the desired privacy guarantees. Additionally, we provide a more general method supporting multiple noise mechanisms, and subsampling methods.
@@
Contributions. (1) We introduce a general framework for differentially private SPRTs, called DP-SPRT, that supports various noise mechanisms while providing privacy guarantees and error control. We propose a generic sample complexity analysis of DP-SPRT for Bernoulli observations and particularize it to Gaussian noise (for Renyi DP) and Laplace noise (for $\varepsilon$-DP). (2) This framework is based on the novel OutsideInterval mechanism for differentially privately comparing a sequence of queries to two thresholds. This might be of independent interest for applications for DP problems with or without continual observations. (3) We further prove a lower bound on the sample complexity of any $\varepsilon$-DP algorithm that achieves prescribed type I and type II errors. This lower bound reveals two distinct regimes: one where privacy has minimal impact on sample complexity, and another where higher privacy requirements substantially increase the expected sample size. It also shows that DP-SPRT with Laplace noise is near optimal in a regime of small errors and when the two hypotheses are close. (4) On the practical side, we use subsampling to propose a practical improvement of DP-SPRT with better privacy-utility trade-off, especially in high privacy regimes.
@@ 2 Preliminaries: Sequential Tests, Differential Privacy, and SPRT
For one-parameter canonical exponential families (Brown,, 1986), in which $\Theta\subseteq\mathbb{R}$ and the density can be written $f_{\theta}(x)=\exp(\theta x-b(\theta))$ where $b:\Theta\rightarrow\mathbb{R}$ is called the log-partition function, SPRT takes a more explicit form (see Appendix A). To fix the ideas, we assume in sequel that $\theta_{0}<\theta_{1}$. Letting $\mu_{i}$ be the mean of the distribution $\nu_{\theta_{i}}$ and $\mathrm{KL}$ the Kullback-Leibler divergence, we have $\tau:=\min(\tau_{0},\tau_{1})$, where
[…]
@@ 3 OutsideInterval Mechanism
(ii) If the noise-adding mechanism corresponding to $\mathcal{D}_{Z}$ has an RDP profile $\varepsilon_{Z}(\alpha)$ for queries with sensitivity $\Delta$ and the one for $\mathcal{D}_{Y}$ has an RDP profile $\varepsilon_{Y}(\alpha)$ for queries with sensitivity $2\Delta$, then OutsideInterval (denoted by $\mathcal{A}$) satisfies:
@@ 4 Differentially Private Sequential Probability Ratio Test
$\tau_{0}$ $:=\inf\bigg{\{}n\in\mathbb{N}_{>0}:\ \frac{\sum_{i=1}^{n}X_{i}}{n}+\frac{Y_{n}}{n}\leq\mu_{0}+\frac{\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}})-\frac{1}{n}\log\left(\frac{1}{\gamma\beta}\right)}{\theta_{1}-\theta_{0}}-\tfrac{1}{n}Z-C\left(n,(1-\gamma)\beta\right)\bigg{\}},$
$\tau_{1}$ $:=\inf\bigg{\{}n\in\mathbb{N}_{>0}:\ \frac{\sum_{i=1}^{n}X_{i}}{n}+\frac{Y_{n}}{n}\geq\mu_{1}-\frac{\mathrm{KL}(\nu_{\theta_{1}},\nu_{\theta_{0}})-\frac{1}{n}\log\left(\frac{1}{\gamma\alpha}\right)}{\theta_{1}-\theta_{0}}+\tfrac{1}{n}Z+C\left(n,(1-\gamma)\alpha\right)\bigg{\}}$
where $Y_{n}\overset{i.i.d.}{\sim}\mathcal{D}_{Y}$ are the successive query noise and $Z\sim\mathcal{D}_{Z}$ is the threshold noise. The decision rule is defined as $\hat{d}=i$ if $\tau=\tau_{i}$. DP-SPRT is described formally in Algorithm 2, where the blue terms highlight the differences with the vanilla SPRT.
@@
8:  $\hat{T}_{0}^{n}\leftarrow\mu_{0}+\frac{\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}})-\log(1/({\color[rgb]{0.3,0.3,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.3,0.3,1}\gamma}\beta))/n}{\theta_{1}-\theta_{0}}{\color[rgb]{0.3,0.3,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.3,0.3,1}-C\left(n,(1-\gamma)\beta\right)}$
9:  $\hat{T}_{1}^{n}\leftarrow\mu_{1}-\frac{\mathrm{KL}(\nu_{\theta_{1}},\nu_{\theta_{0}})-\log(1/({\color[rgb]{0.3,0.3,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.3,0.3,1}\gamma}\alpha))/n}{\theta_{1}-\theta_{0}}{\color[rgb]{0.3,0.3,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.3,0.3,1}+C\left(n,(1-\gamma)\alpha\right)}$
@@
(ii) $\bm{(\alpha,\varepsilon)}$-RDP: If the noise-adding mechanism corresponding to distribution $\mathcal{D}_{Z}$ and $\mathcal{D}_{Z}$ have an RDP profile $\varepsilon_{Z}(\alpha)$ for queries with sensitivity $1$ and $\varepsilon_{Y}(\alpha)$ for queries with sensitivity $2$, then DP-SPRT satisfies:
@@
$\forall\delta\in(0,1),\ \sum_{n=1}^{\infty}\mathbb{P}\left(\frac{Y_{n}}{n}-\frac{Z}{n}>C\left(n,\delta\right)\right)\leq\delta\;.$ (3)
The main idea is to decompose the error probabilities using a union bound into the error from the standard SPRT (bounded via a change-of-measure argument) and the error from privacy noise (controlled through the correction function). The parameter $\gamma$ allocates the total error budget between these two sources. See Appendix D for the full proof.
@@
Let us assume (3) holds. For any error allocation $\gamma\in(0,1)$ and $i\in\{0,1\}$, the expected stopping time of DP-SPRT satisfies:
$\mathbb{E}_{\theta_{0}}[\tau]$ $\leq$ $1+(1-\gamma)\beta+\frac{1}{1-\exp\left(-\frac{(\mathrm{TV}(\nu_{\theta_{0}},\nu_{\theta_{1}}))^{4}}{2(\theta_{1}-\theta_{0})^{2}}\right)}+N(\theta_{0},\theta_{1},\beta,\gamma)$
$\mathbb{E}_{\theta_{1}}[\tau]$ $\leq$ $1+(1-\gamma)\alpha+\frac{1}{1-\exp\left(-\frac{(\mathrm{TV}(\nu_{\theta_{0}},\nu_{\theta_{1}}))^{4}}{2(\theta_{1}-\theta_{0})^{2}}\right)}+N(\theta_{1},\theta_{0},\alpha,\gamma)$
where
$N(\theta,\theta^{\prime},\delta,\gamma)=\inf\left\{n:\frac{\log(1/(\delta\gamma))/n}{|\theta-\theta^{\prime}|}+2C(n,(1-\gamma)\delta)\leq\frac{1}{2}\frac{\mathrm{KL}(\nu_{\theta},\nu_{\theta^{\prime}})}{|\theta-\theta^{\prime}|}\right\}\;.$
@@
Variants of DP-SPRT. The generic DP-SPRT framework provides flexibility in selecting noise distributions based on privacy requirements. Table 1 summarizes two key implementations: Laplace noise for pure $\varepsilon$-DP and Gaussian noise for Rényi DP.
Table 1: Comparison of DP-SPRT instantiations with different noise distributions
Laplace Noise Gaussian Noise
Noise distributions $Y_{n}\sim\text{Lap}\left(\frac{4}{\varepsilon}\right)$ $Y_{n}\sim\mathcal{N}(0,\sigma_{Y}^{2})$
$Z\sim\text{Lap}\left(\frac{2}{\varepsilon}\right)$ $Z\sim\mathcal{N}(0,\sigma_{Z}^{2})$
Correction function $\frac{6\log(n^{s}\zeta(s)/\delta)}{n\varepsilon}$ $\frac{\sqrt{2(\sigma_{Y}^{2}+\sigma_{Z}^{2})\log(n^{s}\zeta(s)/2\delta)}}{n}$
Privacy guarantee $\varepsilon$-Differential Privacy $(\alpha,\varepsilon(\alpha))$-RDP where
$\varepsilon(\alpha)=\frac{\alpha-1/2}{\alpha-1}\cdot\frac{\alpha}{\sigma_{Z}^{2}}+\frac{\alpha}{2\sigma_{Y}^{2}}+\frac{\log\left(2\mathbb{E}_{\mathcal{D}}[\mathbb{E}_{\mathcal{A}(D^{\prime})}[\tau|Z=z]^{2}]\right)}{2(\alpha-1)}$
Both variants satisfy the condition (3), ensuring correctness with the specified error probabilities (see Appendix (C)). The Laplace noise variant provides the strongest privacy guarantee (pure $\varepsilon$-DP), while the Gaussian implementation offers Rényi differential privacy that can be converted to approximate $(\varepsilon,\delta)$-DP using standard techniques (Mironov,, 2017). Detailed sample complexity analyses for these implementations are presented in Appendix E.

### v2

@@ Abstract
We revisit Wald’s celebrated Sequential Probability Ratio Test for sequential tests of two simple hypotheses, under privacy constraints. We propose DP-SPRT, a wrapper that can be calibrated to achieve desired error probabilities and privacy constraints, addressing a significant gap in previous work. DP-SPRT relies on a private mechanism that processes a sequence of queries and stops after privately determining when the query results fall outside a predefined interval. This OutsideInterval mechanism improves upon naive composition of existing techniques like AboveThreshold, achieving a factor-of-2 privacy improvement and thus potentially benefiting other continual monitoring procedures. We prove generic upper bounds on the error and sample complexity of DP-SPRT that can accommodate various noise distributions based on the practitioner’s privacy needs. We exemplify them in two settings: Laplace noise (pure Differential Privacy) and Gaussian noise (Rényi differential privacy). In the former setting, by providing a lower bound on the sample complexity of any $\varepsilon$-DP test with prescribed type I and type II errors, we show that DP-SPRT is near optimal when both errors are small and the two hypotheses are close. Moreover, we conduct an experimental study revealing its good practical performance.
@@ 1 Introduction
Related Work. Differential privacy (DP) (Dwork et al.,, 2006) has emerged as the standard framework for addressing privacy concerns in statistical analyses. By ensuring that the presence or absence of any single data point cannot significantly affect the outcome, DP offers a principled method to protect individual data while enabling meaningful statistical inference. Researchers have explored DP hypothesis testing– primarily in the fixed-sample setting. For example, Gaboardi et al., (2016) studies private chi-squared testing, Canonne et al., (2019) investigates optimal private tests for simple hypotheses, Couch et al., (2019) studies nonparametric hypothesis testing with DP, and Kazan et al., (2023) proposes a generic framework for any fixed-sample hypothesis test. These approaches, however, are not directly applicable to sequential testing due to the adaptive nature of sample collection.
For sequential settings, existing differentially private mechanisms like AboveThreshold (Dwork et al.,, 2006) and the sparse vector technique enable monitoring whether query results exceed specified thresholds while preserving privacy. However, many applications—including sequential hypothesis testing—require monitoring both upper and lower bounds simultaneously, leading to suboptimal privacy guarantees when using naive composition.
Wang et al., (2022) explored the privatization of SPRT by adding Laplace noise to thresholds and using the exponential mechanism for the final decision. However, due to their specific use case, their approach provides a weaker notion of privacy that only requires that critical privacy failures do not happen only for a subset of plausible datasets. In contrast, our work maintains DP throughout the testing process while working to minimize stopping time. Zhang et al., (2022) developed PrivSPRT, a private version of the SPRT using Gaussian noise with Renyi DP guarantees (Mironov,, 2017). The authors provide upper bounds on the error probabilities given fixed thresholds and noise level, but do not specify how to calibrate these thresholds to achieve desired Type I and Type II error rates while maintaining a desired privacy guarantee. In practice, this forces to rely on Monte Carlo simulation to determine appropriate thresholds, a process that may itself leak privacy in certain scenarios (Papernot and Steinke,, 2021).
@@
Contributions. (1) We introduce a general framework for differentially private SPRTs, called DP-SPRT, that supports various noise mechanisms while providing privacy guarantees and error control. We propose a generic sample complexity analysis of DP-SPRT for Bernoulli observations and particularize it to Gaussian noise (for Renyi DP) and Laplace noise (for $\varepsilon$-DP). (2) This framework is based on the novel OutsideInterval mechanism for differentially privately comparing a sequence of queries to two thresholds. This might be of independent interest for applications to DP problems with or without continual observations. (3) We further prove a lower bound on the sample complexity of any $\varepsilon$-DP algorithm that achieves prescribed type I and type II errors. This lower bound reveals two distinct regimes: one where privacy has minimal impact on sample complexity, and another where higher privacy requirements substantially increase the expected sample size. It also shows that DP-SPRT with Laplace noise is near optimal in a regime of small errors and when the two hypotheses are close. (4) On the practical side, we use subsampling to propose a practical improvement of DP-SPRT with better privacy-utility trade-off, especially in high privacy regimes.
@@ 2 Preliminaries: Sequential Tests, Differential Privacy, and SPRT
For one-parameter canonical exponential families (Brown,, 1986), in which $\Theta\subseteq\mathbb{R}$ and the density can be written $f_{\theta}(x)=\exp(\theta x-b(\theta))$ where $b:\Theta\rightarrow\mathbb{R}$ is called the log-partition function, SPRT takes a more explicit form (see Appendix A). To fix the ideas, we assume in sequel that $\theta_{0}<\theta_{1}$. Letting $\bar{X}_{n}=\sum_{i=1}^{n}X_{i}/n$ denote the empirical mean, $\mu_{i}$ be the mean of the distribution $\nu_{\theta_{i}}$, and $\mathrm{KL}$ the Kullback-Leibler divergence, we have $\tau:=\min(\tau_{0},\tau_{1})$, where
$\begin{split}\tau_{0}&:=\inf\bigg\{n\in\mathbb{N}_{>0}:\bar{X}_{n}\leq\ell_{0}^{n}\bigg\},\\ \tau_{1}&:=\inf\bigg\{n\in\mathbb{N}_{>0}:\bar{X}_{n}\geq\ell_{1}^{n}\bigg\}\end{split}$ (1)
with thresholds:
$\ell_{0}^{n}$ $=\mu_{0}+\frac{\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}})-\log(1/\beta)/n}{\theta_{1}-\theta_{0}},$
$\ell_{1}^{n}$ $=\mu_{1}-\frac{\mathrm{KL}(\nu_{\theta_{1}},\nu_{\theta_{0}})-\log(1/\alpha)/n}{\theta_{1}-\theta_{0}}.$
@@ 3 OutsideInterval Mechanism
(ii) If the noise-adding mechanism corresponding to $\mathcal{D}_{Z}$ has an RDP profile $\varepsilon_{Z}(\alpha)$ for queries with sensitivity $\Delta$ and the one for $\mathcal{D}_{Y}$ has an RDP profile $\varepsilon_{Y}(\alpha)$ for queries with sensitivity $2\Delta$, then OutsideInterval (denoted by $\mathcal{A}$) satisfies: for all $\alpha>1$,
@@ 4.1 General Definition
$\tau_{0}$ $:=\inf\bigg\{n\in\mathbb{N}_{>0}:\bar{X}_{n}+\frac{Y_{n}}{n}\leq\tilde{\ell}_{0}^{n}-\frac{Z}{n}\bigg\},$
$\tau_{1}$ $:=\inf\bigg\{n\in\mathbb{N}_{>0}:\bar{X}_{n}+\frac{Y_{n}}{n}\geq\tilde{\ell}_{1}^{n}+\frac{Z}{n}\bigg\}$
where the private thresholds are:
$\tilde{\ell}_{0}^{n}$ $=\mu_{0}+\frac{\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}})-\frac{\log(1/(\gamma\beta))}{n}}{\theta_{1}-\theta_{0}}-C(n,(1-\gamma)\beta),$
$\tilde{\ell}_{1}^{n}$ $=\mu_{1}-\frac{\mathrm{KL}(\nu_{\theta_{1}},\nu_{\theta_{0}})-\frac{\log(1/(\gamma\alpha))}{n}}{\theta_{1}-\theta_{0}}+C(n,(1-\gamma)\alpha).$
Here, $Y_{n}\overset{i.i.d.}{\sim}\mathcal{D}_{Y}$ are the successive query noise and $Z\sim\mathcal{D}_{Z}$ is the threshold noise. The decision rule is defined as $\hat{d}=i$ if $\tau=\tau_{i}$. DP-SPRT is described formally in Algorithm 2, where the blue terms highlight the differences with the vanilla SPRT.
@@
8:  $\hat{T}_{0}^{n}\!\leftarrow\mu_{0}+\frac{\mathrm{KL}(\nu_{\theta_{0}},\nu_{\theta_{1}})+\log({\color[rgb]{0.3,0.3,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.3,0.3,1}\gamma}\beta)/n}{\theta_{1}-\theta_{0}}{\color[rgb]{0.3,0.3,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.3,0.3,1}-C\left(n,(1-\gamma)\beta\right)}$
9:  $\hat{T}_{1}^{n}\!\leftarrow\mu_{1}-\frac{\mathrm{KL}(\nu_{\theta_{1}},\nu_{\theta_{0}})+\log({\color[rgb]{0.3,0.3,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.3,0.3,1}\gamma}\alpha)/n}{\theta_{1}-\theta_{0}}{\color[rgb]{0.3,0.3,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.3,0.3,1}+C\left(n,(1-\gamma)\alpha\right)}$
@@ 4.2 Generic analysis
(ii) $\bm{(\alpha,\varepsilon)}$-RDP: If the noise-adding mechanism corresponding to distribution $\mathcal{D}_{Z}$ and $\mathcal{D}_{Y}$ have an RDP profile $\varepsilon_{Z}(\alpha)$ for queries with sensitivity $1$ and $\varepsilon_{Y}(\alpha)$ for queries with sensitivity $2$, then DP-SPRT satisfies: for all $\alpha>1$,
@@
$\forall\delta\in(0,1),$ $\ \sum_{n=1}^{\infty}\mathbb{P}\left(\frac{Y_{n}}{n}-\frac{Z}{n}>C\left(n,\delta\right)\right)\leq\delta\;,$ (3)
$\forall\delta\in(0,1),$ $\ \sum_{n=1}^{\infty}\mathbb{P}\left(\frac{Y_{n}}{n}+\frac{Z}{n}<-C\left(n,\delta\right)\right)\leq\delta\;.$ (4)
In particular, if the distribution $\mathcal{D}_{Y}$ is symmetric around zero, we remark that condition (3) alone is sufficient, as it implies condition (4).
The main idea is to decompose the error probabilities using a union bound into the error from the standard SPRT (bounded via a change-of-measure argument) and the error from privacy noise (controlled through the correction function). The parameter $\gamma$ allocates the total error budget between these two sources. The symmetry condition ensures that both tail bounds needed for Type I and Type II error control are satisfied simultaneously. See Appendix D for the full proof.
@@
Let us assume conditions (3) and (4) hold. For any error allocation $\gamma\in(0,1)$, the expected stopping time of DP-SPRT satisfies:
$\mathbb{E}_{\theta_{0}}[\tau]$ $\leq 1+(1-\gamma)\beta+\frac{1}{1-e^{-\frac{(\mathrm{TV}(\nu_{\theta_{0}},\nu_{\theta_{1}}))^{4}}{2(\theta_{1}-\theta_{0})^{2}}}}+N_{0},$
$\mathbb{E}_{\theta_{1}}[\tau]$ $\leq 1+(1-\gamma)\alpha+\frac{1}{1-e^{-\frac{(\mathrm{TV}(\nu_{\theta_{0}},\nu_{\theta_{1}}))^{4}}{2(\theta_{1}-\theta_{0})^{2}}}}+N_{1}$
where $N_{0}=N(\theta_{0},\theta_{1},\beta,\gamma)$, $N_{1}=N(\theta_{1},\theta_{0},\alpha,\gamma)$, and
$N(\theta,\theta^{\prime},\delta,\gamma)=\inf\bigg\{n:\frac{\log(1/(\delta\gamma))}{n|\theta-\theta^{\prime}|}+2C(n,(1-\gamma)\delta)\leq\frac{\mathrm{KL}(\nu_{\theta},\nu_{\theta^{\prime}})}{2|\theta-\theta^{\prime}|}\bigg\}.$
@@
4.3 Variants of DP-SPRT
The generic DP-SPRT framework provides flexibility in selecting noise distributions based on privacy requirements. Table 1 summarizes two key implementations: Laplace noise for pure $\varepsilon$-DP and Gaussian noise for Rényi DP.
Table 1: Comparison of DP-SPRT instantiations
Laplace Gaussian
Query noise $Y_{n}\sim\text{Lap}(4/\varepsilon)$ $Y_{n}\sim\mathcal{N}(0,\sigma_{Y}^{2})$
Threshold noise $Z\sim\text{Lap}(2/\varepsilon)$ $Z\sim\mathcal{N}(0,\sigma_{Z}^{2})$
Correction $\frac{6\log(n^{s}\zeta(s)/\delta)}{n\varepsilon}$ $\frac{\sqrt{2\sigma^{2}\log(n^{s}\zeta(s)/2\delta)}}{n}$
Privacy $\varepsilon$-DP $(\alpha,\varepsilon(\alpha))$-RDP
Here, $\sigma^{2}=\sigma_{Y}^{2}+\sigma_{Z}^{2}$ and $\varepsilon(\alpha)=\frac{\alpha-1/2}{\alpha-1}\cdot\frac{\alpha}{\sigma_{Z}^{2}}+\frac{\alpha}{2\sigma_{Y}^{2}}+\frac{\log(2\mathbb{E}[\mathbb{E}[\tau|Z]^{2}])}{2(\alpha-1)}$.
Both variants satisfy conditions (3) and (4), ensuring correctness with the specified error probabilities. The Laplace noise variant provides the strongest privacy guarantee (pure $\varepsilon$-DP), while the Gaussian implementation offers Rényi differential privacy that can be converted to approximate $(\varepsilon,\delta)$-DP using standard techniques (Mironov,, 2017). Detailed sample complexity analyses for these implementations are presented in Appendix E.